# VideoRoPE: What Makes for Good Video Rotary Position Embedding?

### v2

@@ Abstract
While Rotary Position Embedding (RoPE) and its variants are widely adopted for their long-context capabilities, the extension of the 1D RoPE to video, with its complex spatio-temporal structure, remains an open challenge. This work first introduces a comprehensive analysis that identifies four key characteristics essential for the effective adaptation of RoPE to video, which have not been fully considered in prior work. As part of our analysis, we introduce a challenging V-NIAH-D (Visual Needle-In-A-Haystack with Distractors) task, which adds periodic distractors into V-NIAH. The V-NIAH-D task demonstrates that previous RoPE variants, lacking appropriate temporal dimension allocation, are easily misled by distractors. Based on our analysis, we introduce VideoRoPE, with a 3D structure designed to preserve spatio-temporal relationships. VideoRoPE features low-frequency temporal allocation to mitigate periodic oscillations, a diagonal layout to maintain spatial symmetry, and adjustable temporal spacing to decouple temporal and spatial indexing. VideoRoPE consistently surpasses previous RoPE variants, across diverse downstream tasks such as long video retrieval, video understanding, and video hallucination. Our code will be available at https://github.com/Wiselnn570/VideoRoPE.
@@ 1 Introduction
Figure 2: Left: To demonstrate the importance of frequential allocation, based on VIAH (a) we present a more challenging V-NIAH-D task (b) that similar images are inserted as distractors. Right: Compared to M-RoPE, our VideoRoPE is more robust in retrieval and is less affected by distractors.
@@ 4 VideoRoPE
Low-frequency Temporal Allocation (LTA). As shown in Eq. (2), the vanilla RoPE (Su et al., 2024) uses all dimensions to model the 1D position information. And as indicated in Eq. (4), M-RoPE (Wang et al., 2024a) uses dimensions to model temporal, horizontal, and vertical dimensions sequentially. However, previous frequency allocation strategies are suboptimal because different RoPE dimensions capture dependencies at varying ranges. As shown in Fig. 3, an interesting observation is that the local attention branch (as reported in (Han et al., 2024)) corresponds to lower dimensions, while the global branch (or attention sink, as in (Xiao et al., 2023)) corresponds to higher dimensions. To sum up, lower dimensions (higher frequency, shorter monotonic intervals, larger $\theta_{n}$) tend to capture relative distances and local semantics (Men et al., 2024; Barbero et al., 2024), while higher dimensions (lower frequency, wider monotonic intervals, smaller $\theta_{n}$) capture longer-range dependencies (Barbero et al., 2024).
@@
As shown in Fig. 4 (a), M-RoPE’s temporal position embeddings are significantly distorted by periodic oscillations (Men et al., 2024), leading to identical embeddings for distant positions. For instance, considering the last three rotary angles, the temporal embeddings are severely affected by these oscillations due to their short monotonic intervals (and even shorter intervals in lower dimensions). This periodicity creates “hash collisions” (red planes), where distant positions share near-identical embeddings, making the model susceptible to distractor influence. Fortunately, our VideoRoPE (Fig. 4 (b)) is free from oscillation and Hash collision in temporal modeling. The visualized relationship between the periodicity, monotonicity, and temporal modeling.
@@ 5.1 Experimental Setup
Implementation Details. Using the aforementioned video training data, we fine-tune different modes that use different positional encoding strategies, such as the Vanilla RoPE (Su et al., 2024), Time-Aware Dual RoPE (TAD-RoPE) (Gao et al., 2024), M-RoPE (Wang et al., 2024a), and our VideoRoPE. All models are initialized with the Vision Transformer from Qwen2-VL-7B and LLM (Vanilla RoPE) from Qwen2-7B (Yang et al., 2024a). Our fine-tuning incorporates our VideoRoPE to process the spatiotemporal nature of the video data effectively. We adopt Qwen2-VL’s fine-tuning settings, processing each video at 2 fps with a maximum of 128 frames and dynamically adjusting the image resolution to maintain a consistent token count. However, to prevent memory overflow, we use a context window of 8192 tokens.
@@ 5.5 Ablation Studies
Ablation Studies on Module Design. We conduct a series of ablation experiments on the modules introduced in Section 4. Through these experiments, we quantitatively evaluate the performance improvements, specifically analyzing their impact on the LongVideoBench and MLVU benchmarks. The experimental results are presented in Tab. 5. The baseline setting, M-RoPE (Wang et al., 2024a), achieves scores of 54.35 on LongVideoBench and 61.10 on MLVU (both using a 64k context length). By progressively integrating the DL (Diagonal Layout), LTA (Low-frequency Temporal Allocation), and ATS (Adjustable Temporal Spacing) modules, our method shows a continuous improvement in performance, achieving enhanced scores of 57.26 on LongVideoBench and 65.56 on MLVU (both using a 64k context length). These results demonstrate the effectiveness of our approach in leveraging spatial-temporal positional information. To refine the discussion on the allocation of $x$ and $y$ within LTA, we quantitatively examine the performance impact of interleaved versus sequential allocations. Additionally, we explore the optimal scaling factor in ATS by adjusting its values. Please refer to Appendix A.1 for more ablation studies on the scaling factor $\delta$ and the $x$/$y$ interleaved allocation.
@@ A.1 Supplementary Ablation Experiments
Ablation Studies on the Scaling Factor $\bm{\delta}$ for ATS. This section extends the discussion from Section 4, further examining how varying temporal information—achieved by adjusting the scaling factor $\delta$—affects the alignment between text and video tokens. The proper alignment of these tokens is key for improving the model’s understanding of both temporal and semantic relationships between text and video. As shown in Table 6, we conduct evaluations using the LongVideoBench benchmark, which provides a comprehensive evaluation of video understanding tasks. By sampling $\delta$ values in the range $[0.5,3]$ with an interval of $0.5$, we observe that performance improves with increasing $\delta$ up to a point. Our findings indicate that the optimal performance is achieved when $\delta=2$, suggesting that this scaling factor best balances the temporal alignment between text and video tokens.
Table 6: Ablation Study About Different Scaling Factor $\delta$.
Scaling Factor $\bm{\delta}$ LongVideoBench Average
8k 16k 32k 64k
0.5 51.92 53.52 52.80 52.07 52.57
1.0 54.46 55.49 54.66 55.60 55.05
1.5 54.35 55.00 55.31 55.91 55.14
2.0 54.46 55.29 57.15 57.26 56.04
2.5 53.42 53.73 54.25 55.08 54.12
3.0 53.63 53.63 53.63 55.18 54.01

### v3

@@ Abstract
While Rotary Position Embedding (RoPE) and its variants are widely adopted for their long-context capabilities, the extension of the 1D RoPE to video, with its complex spatio-temporal structure, remains an open challenge. This work first introduces a comprehensive analysis that identifies four key characteristics essential for the effective adaptation of RoPE to video, which have not been fully considered in prior work. As part of our analysis, we introduce a challenging V-NIAH-D (Visual Needle-In-A-Haystack with Distractors) task, which adds periodic distractors into V-NIAH. The V-NIAH-D task demonstrates that previous RoPE variants, lacking appropriate temporal dimension allocation, are easily misled by distractors. Based on our analysis, we introduce VideoRoPE, with a 3D structure designed to preserve spatio-temporal relationships. VideoRoPE features low-frequency temporal allocation to mitigate periodic oscillations, a diagonal layout to maintain spatial symmetry, and adjustable temporal spacing to decouple temporal and spatial indexing. VideoRoPE consistently surpasses previous RoPE variants, across diverse downstream tasks such as long video retrieval, video understanding, and video hallucination. Our code is available at https://github.com/Wiselnn570/VideoRoPE.
@@ 1 Introduction
Figure 2: Left: To demonstrate the importance of frequential allocation, based on VIAH (a) we present a more challenging V-NIAH-D task (b) that similar images are inserted as distractors. Right: Compared to M-RoPE, our VideoRoPE is more robust in retrieval and is less affected by distractors. See Fig. 7 in the Experiments section for details on the horizontal and vertical axes.
@@ 4 VideoRoPE
Low-frequency Temporal Allocation (LTA). As shown in Eq. (2), the vanilla RoPE (Su et al., 2024) uses all dimensions to model the 1D position information. And as indicated in Eq. (4), M-RoPE (Wang et al., 2024a) uses different dimensions to model temporal, horizontal, and vertical dimensions sequentially. However, previous frequency allocation strategies are suboptimal because different RoPE dimensions capture dependencies at varying ranges. As shown in Fig. 3, an interesting observation is that the local attention branch (as reported in (Han et al., 2024)) corresponds to lower dimensions, while the global branch (or attention sink, as in (Xiao et al., 2023)) corresponds to higher dimensions. To sum up, lower dimensions (higher frequency, shorter monotonic intervals, larger $\theta_{n}$) tend to capture relative distances and local semantics (Men et al., 2024; Barbero et al., 2024), while higher dimensions (lower frequency, wider monotonic intervals, smaller $\theta_{n}$) capture longer-range dependencies (Barbero et al., 2024).
@@
As shown in Fig. 4 (a), M-RoPE’s temporal position embeddings are significantly distorted by periodic oscillations (Men et al., 2024), leading to identical embeddings for distant positions. For instance, considering the last three rotary angles, the temporal embeddings are severely affected by these oscillations due to their short monotonic intervals (and even shorter intervals in lower dimensions). This periodicity creates “hash collisions” (red planes), where distant positions share near-identical embeddings, making the model susceptible to distractor influence. Fortunately, our VideoRoPE (Fig. 4 (b)) is free from oscillation and Hash collision in temporal modeling. The relationship between periodicity, monotonicity, and temporal modeling is visualized in Fig 4.
@@ 5.1 Experimental Setup
Implementation Details. Using the aforementioned video training data, we fine-tune different models that use different positional encoding strategies, such as the Vanilla RoPE (Su et al., 2024), Time-Aware Dual RoPE (TAD-RoPE) (Gao et al., 2024), M-RoPE (Wang et al., 2024a), and our VideoRoPE. All models are initialized with the Vision Transformer from Qwen2-VL-7B and LLM (Vanilla RoPE) from Qwen2-7B (Yang et al., 2024a). Our fine-tuning incorporates our VideoRoPE to process the spatiotemporal nature of the video data effectively. We adopt Qwen2-VL’s fine-tuning settings, processing each video at 2 fps with a maximum of 128 frames and dynamically adjusting the image resolution to maintain a consistent token count. However, to prevent memory overflow, we use a context window of 8192 tokens.
@@ 5.5 Ablation Studies
Ablation Studies on Module Design.
We conduct ablation experiments on the modules introduced in Section 4, quantitatively evaluating their impact on LongVideoBench and MLVU benchmarks. The experimental results are presented in Tab. 5. The baseline setting, M-RoPE (Wang et al., 2024a), achieves scores of 54.35 on LongVideoBench and 61.10 on MLVU (both using a 64k context length). By progressively integrating the DL (Diagonal Layout), LTA (Low-frequency Temporal Allocation), and ATS (Adjustable Temporal Spacing) modules, our method shows a continuous improvement in performance, achieving enhanced scores of 57.26 on LongVideoBench and 65.56 on MLVU (both using a 64k context length). These results demonstrate the effectiveness of our approach in leveraging spatial-temporal positional information. To refine the analysis of $x$ and $y$ allocation in LTA, we quantitatively evaluate interleaved vs. sequential layouts. We also compare strategies for allocating $t$, $x$, and $y$, including M-RoPE, a uniform interleaved layout, and our VideoRoPE design. Additionally, we explore the optimal ATS scaling factor by varying its value, and further ablate the diagonal layout module to validate the symmetry-based design. See Appendix A.1 for details.
@@ 6 Conclusion
Acknowledgments
This work was supported by National Key R&D Program of China 2022ZD0161600, Shanghai Artificial lntelligence Laboratory, the Centre for Perceptual and Interactive Intelligence (CPII) Ltd under the Innovation and Technology Commission (ITC)’s InnoHK. Dahua Lin is a PI of CPII under the InnoHK.
@@ A.1 Supplementary Ablation Experiments
Ablation Studies on the Scaling Factor $\bm{\delta}$ for ATS. We conduct a series of experiments to further investigate the impact of the temporal scaling factor $\delta$ on the alignment between video and text representations. Accurate temporal alignment plays a vital role in enhancing the model’s understanding of both semantic and sequential aspects of video-language data. To this end, we evaluate the model’s performance across three representative video-language benchmarks—LongVideoBench, MLVU, and VideoMME—by varying the temporal scaling factor $\delta$ from 0.5 to 3.0. As shown in Table 6, we observe a consistent trend across all benchmarks: performance improves as $\delta$ increases, peaking at $\delta=2$ with an average score of 60.92. These results suggest that setting $\delta=2$ strikes the best balance between temporal resolution and semantic alignment, resulting in optimal overall performance.
Table 6: Performance under different scaling factors $\delta$ across multiple benchmarks.
Scaling Factor $\bm{\delta}$ LongVideoBench MLVU VideoMME Avg
0.5 50.83 59.87 58.33 56.34
1.0 54.11 63.54 59.67 59.11
2.0 55.50 65.59 61.67 60.92
3.0 53.83 63.38 60.33 59.18
@@
Ablation Studies on Diagonal Layout Validated on More Benchmarks
To further substantiate our claim that the Diagonal Layout (DL) enhances a model’s capability in video understanding tasks, we conduct additional ablation studies on four diverse and challenging benchmarks: MLVU, VideoHallucer, V-NIAH, and V-NIAH-D. These benchmarks cover a wide range of evaluation perspectives, from multi-level video question answering to hallucination detection and fine-grained temporal alignment. As shown in Table 8, incorporating the DL module consistently improves performance over the baseline model across all benchmarks. Specifically, we observe notable gains on MLVU and V-NIAH and V-NIAH-D, suggesting that DL effectively facilitates better temporal reasoning and semantic alignment. These results reinforce the generalizability and robustness of the proposed Diagonal Layout design in understanding across various tasks.
Table 8: Effect of Diagonal Layout (DL) across multiple benchmarks.
Method MLVU VideoHallucer V-NIAH V-NIAH-D
baseline 61.56 34.3 78.67 74.67
+ DL 63.03 34.8 80.44 76.44
Ablation Studies on Different Frequency Allocation Strategies
We compare three different frequency allocation strategies: the M-RoPE approach, which emphasizes high-frequency modeling of temporal information and follows a [t…x…y…] format; an interleaved and evenly distributed pattern such as [t t x y x y x y]; and our proposed VideoRoPE method, which prioritizes positional encoding followed by low-frequency temporal modeling, arranged in a [x y…t…] format.
We evaluate these approaches on the LongVideoBench benchmark under varying context lengths. This benchmark includes a diverse set of video scenarios, ranging from rapidly changing dynamic scenes to slowly evolving static content.
As shown in the results below, our low-frequency temporal allocation consistently outperforms the interleaved [t t x y x y x y] pattern on average. This suggests that our frequency design more effectively balances global temporal context modeling with local spatial dynamics, making it better suited to handle a wide variety of video conditions.
Table 9: Comparison of different frequency allocation strategies under various context lengths.
Context [t…x…y…] [t t x y x y] [xy…t…] (Ours)
16k 60.05 59.95 62.03
32k 59.33 58.40 59.54
64k 58.71 57.73 59.12
Avg 59.36 59.06 60.14